# Mean-field Underdamped Langevin Dynamics and its Space-Time Discretization

### 26 Dec 2023 v1

@@ Appendix D Discretization analysis
In this section, we provide the proof of Theorem 3 and Theorem 4 establishing the global convergence of the discrete-time-space processes. Our discretization analysis is unified for the MULA and NULA.
D.1 Proof of Theorem 3
@@
where the first inequality follows from Assumption 2 and the last inequality follows from Lemma 6 and the inequality $\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}\right)^{2}\leq\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}$:
@@
$\textsf{KL}(\mathbf{Q}_{Nh}\|\mathbf{p}_{Nh})$ $\leq\frac{16\mathscr{L}^{4}h^{5}}{\gamma}\sum_{k=0}^{K-1}\mathbb{E}_{\mathbf{Q}_{Kh}}\|x_{kh}\|^{2}+\frac{64\mathscr{L}^{2}h^{3}}{\gamma}\sum_{k=0}^{K-1}\mathbb{E}_{\mathbf{Q}_{Kh}}\|v_{kh}\|^{2}+\frac{16\mathscr{L}^{4}h^{5}K}{\gamma}+{32\mathscr{L}^{2}h^{4}Kd}$
@@
where $T=Kh$. By Lemma 2 and Theorem 1, we obtain
@@ D.2 Proof of Theorem 4
where the first inequality follows from Assumption 2 and the last inequality follows from Lemma 6 and the inequality $\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}\right)^{2}\leq\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}$:
@@
$\frac{1}{N}\sum_{i=1}^{N}\textsf{KL}(\mathbf{Q}^{i}_{Kh}\|\mathbf{P}^{i}_{Kh})$ $\leq\frac{16\mathscr{L}^{4}h^{5}}{\gamma}\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\|x^{i}_{kh}\|^{2}+\frac{64\mathscr{L}^{2}h^{3}}{\gamma}\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\|v_{kh}\|^{2}$
@@
where $T=Kh$. By Lemma 3 and Theorem 2, we obtain
$\frac{1}{N}\textsf{KL}(\mu_{T}^{N}\|\mu_{*}^{\otimes N})\leq\frac{1}{N}\mathcal{F}^{N}(\mu_{T}^{N})-\mathcal{F}(\mu_{*})\leq\left(\frac{1}{N}\mathcal{E}^{N}(\mu_{0}^{N})-\mathcal{E}(\mu_{*})\right)\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{6\sqrt{\mathscr{L}}}T\right)+\frac{60\mathscr{L}d}{N\mathscr{C}_{\textsf{LSI}}}+\frac{36\mathscr{L}^{2}d}{N\mathscr{C}_{\textsf{LSI}}^{2}}$ (42)
Combining with (42), we upper bound the averaged TV distance between $\bar{\mu}_{K}^{i}$ and $\mu_{*}$ over $N$ particles as follows:
@@
$N=\Theta\left(\frac{\mathscr{L}^{2}d}{\mathscr{C}_{\textsf{\em LSI}}^{2}\epsilon^{2}}\right).$ (45)
@@ Appendix E Experimental settings
Parameters Value
$\varphi_{0}$ $10^{-4}$
$\varphi_{1}$ $0.02$
$\varphi_{2}$ $0.99$
$\varphi_{3}$ $0.02$
$\eta$ $10^{-3}$
$h_{1}$ $0.01$
$h_{2}$ $0.01$
$h_{3}$ $0.01$
$\lambda_{1}$ $10^{-4}$
$\lambda_{2}$ $10^{-4}$
@@ Appendix F Methods for comparisons
where $\delta_{hk}\triangleq\mathbb{E}\|{D}_{\mu}F(\bar{\mu}_{k+1},x_{k+1})-{D}_{\mu}F(\bar{\mu}_{k},x_{k})\|^{2}$ and $\mathbb{E}$ is taken under the joint law of $\bar{\mu}_{k+1}$ and $\bar{\mu}_{k}$. Now we bound $\delta_{hk}$ uniformly in $k$ with a different method from the one in Nitanda et al. (2022). We do not need to specify $F$ to be the objective of training nerual networks. Since $F$ is $\mathscr{L}$-smooth and satisfies Assumption 4, we obtain
@@
We refer to Lemma 1 in Suzuki et al. (2023) to uniformly bound $\mathbb{E}\|x_{k}\|^{2}$. Before applying Lemma 1, we translate some constants in Suzuki et al. (2023) into our constants systems. Suzuki et al. (2023) assumes that $\|{D}_{\mu}U(\mu,x)\|\leq R$, $\lambda_{1}I_{d}\preceq\nabla^{2}r(x)\preceq\lambda_{2}I_{d}$. We let $R=\mathscr{L}$ and $\lambda_{2}=\mathscr{L}$ (since this specification matches our Assumption 4). We prove Lemma 1 proposed by Suzuki et al. (2023) in the mean-field setting without particle approximation. But we also assume the decomposition $F(\mu)=U(\mu)+\mathbb{E}_{x\sim\mu}[r(x)]$ with $\|{D}_{\mu}U(\mu,x)\|\leq\mathscr{L}$ and $\lambda_{1}I_{d}\preceq\nabla^{2}r\preceq\mathscr{L}I_{d}$. Given the update of the MLA, if $h\leq\frac{\lambda_{1}}{2\mathscr{L}^{2}}$, we have

### 01 Jan 2024 v2

@@ Appendix D Discretization analysis
In this section, we provide the proof of Theorem 3.3 and Theorem 3.4 establishing the global convergence of the discrete-time-space processes. Our discretization analysis is unified for the MULA and NULA.
D.1 Proof of Theorem 3.3
@@
where the first inequality follows from Assumption 2.2 and the last inequality follows from Lemma 6 and the inequality $\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}\right)^{2}\leq\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}$:
@@
$\textsf{KL}(\mathbf{Q}_{Kh}\|\mathbf{p}_{Kh})$ $\leq\frac{16\mathscr{L}^{4}h^{5}K}{\gamma}\max_{0\leq k\leq K}\mathbb{E}_{\mathbf{Q}_{Kh}}\|x_{kh}\|^{2}+\frac{64\mathscr{L}^{2}h^{3}K}{\gamma}\max_{0\leq k\leq K}\mathbb{E}_{\mathbf{Q}_{Kh}}\|v_{kh}\|^{2}$
$\quad+\frac{16\mathscr{L}^{4}h^{5}K}{\gamma}+{32\mathscr{L}^{2}h^{4}Kd}$
@@
where $T=Kh$. By Lemma 2 and Theorem 3.1, we obtain
@@
where the first inequality follows from Assumption 2.2 and the last inequality follows from Lemma 6 and the inequality $\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}\right)^{2}\leq\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}$:
@@
$\frac{1}{N}\sum_{i=1}^{N}\textsf{KL}(\mathbf{Q}^{i}_{Kh}\|\mathbf{P}^{i}_{Kh})$ $\leq\frac{16\mathscr{L}^{4}h^{5}K}{\gamma}\frac{1}{N}\sum_{i=1}^{N}\max_{0\leq k\leq K}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\|x^{i}_{kh}\|^{2}+\frac{64\mathscr{L}^{2}h^{3}K}{\gamma}\frac{1}{N}\sum_{i=1}^{N}\max_{0\leq k\leq K}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\|v_{kh}\|^{2}$
@@
where $T=Kh$. By Lemma 3 and Theorem 3.2, we obtain
$\frac{1}{N}\textsf{KL}(\mu_{T}^{N}\|\mu_{*}^{\otimes N})\leq\frac{1}{N}\mathcal{F}^{N}(\mu_{T}^{N})-\mathcal{F}(\mu_{*})\leq\frac{\mathcal{E}_{0}^{N}}{N}\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{6\sqrt{\mathscr{L}}}T\right)+\frac{60\mathscr{L}d}{N\mathscr{C}_{\textsf{LSI}}}+\frac{36\mathscr{L}^{2}d}{N\mathscr{C}_{\textsf{LSI}}^{2}},$ (42)
where $\mathcal{E}_{0}^{N}\triangleq\mathcal{E}^{N}(\mu_{0}^{N})-N\mathcal{E}(\mu_{*}).$ Combining with (42), we upper bound the averaged TV distance between $\bar{\mu}_{K}^{i}$ and $\mu_{*}$ over $N$ particles as follows:
@@
$N=\Theta\left(\frac{\mathscr{L}^{2}d}{\mathscr{C}_{\textsf{LSI}}^{2}\epsilon^{2}}\right).$ (45)
@@ Appendix E Experimental settings
Parameters $\varphi_{0}$ $\varphi_{1}$ $\varphi_{2}$ $\varphi_{3}$ $\eta$ $h_{1}$ $h_{2}$ $h_{3}$ $\lambda_{1}$ $\lambda_{2}$
Value $10^{-4}$ $0.02$ $0.99$ $0.02$ $10^{-3}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-4}$ $10^{-4}$
@@ Appendix F Methods for comparisons
where $\delta_{hk}\triangleq\mathbb{E}\|{D}_{\mu}F(\bar{\mu}_{k+1},x_{k+1})-{D}_{\mu}F(\bar{\mu}_{k},x_{k})\|^{2}$ and $\mathbb{E}$ is taken under the joint law of $\bar{\mu}_{k+1}$ and $\bar{\mu}_{k}$. Now we bound $\delta_{hk}$ uniformly in $k$ with a different method from the one in Nitanda et al. (2022). We do not need to specify $F$ to be the objective of training nerual networks. Since $F$ is $\mathscr{L}$-smooth and satisfies Assumption 2.4, we obtain
@@
We refer to Lemma 1 in Suzuki et al. (2023) to uniformly bound $\mathbb{E}\|x_{k}\|^{2}$. Before applying Lemma 1, we translate some constants in Suzuki et al. (2023) into our constants systems. Suzuki et al. (2023) assumes that $\|{D}_{\mu}U(\mu,x)\|\leq R$, $\lambda_{1}I_{d}\preceq\nabla^{2}r(x)\preceq\lambda_{2}I_{d}$. We let $R=\mathscr{L}$ and $\lambda_{2}=\mathscr{L}$ (since this specification matches our Assumption 2.4). We prove Lemma 1 proposed by Suzuki et al. (2023) in the mean-field setting without particle approximation. But we also assume the decomposition $F(\mu)=U(\mu)+\mathbb{E}_{x\sim\mu}[r(x)]$ with $\|{D}_{\mu}U(\mu,x)\|\leq\mathscr{L}$ and $\lambda_{1}I_{d}\preceq\nabla^{2}r\preceq\mathscr{L}I_{d}$. Given the update of the MLA, if $h\leq\frac{\lambda_{1}}{2\mathscr{L}^{2}}$, we have